OpenStudy (anonymous):

find the axis intercepts of the lines y= 4x-7 please help, my exam is tomorrow :(

OpenStudy (anonymous):

When you want the x intercept, you set y = 0, and solve for x. This gives:\[y = 4x-7 \iff 0 = 4x-7 \iff 7 = 4x \iff x = \frac{7}{4}\]If you want the y intercept, set x = 0 and solve for y:\[y = 4x-7 \iff y = 4(0)-7 \iff y = 0-7 \iff y = -7\]

OpenStudy (anonymous):

thankyou so much